OpenStudy (howard-wolowitz):

Can someone explain to me how to do this:

OpenStudy (howard-wolowitz):

each time i try to solve that I get a fraction

OpenStudy (solomonzelman):

those are separate questions, right ?

OpenStudy (solomonzelman):

the one that starts from "Determine the value" and the other that starts from "The graph of a 4\(^{th}\) degree"

OpenStudy (howard-wolowitz):

yes they are, sorry

OpenStudy (solomonzelman):

ok, even if instead of "1" on top of the x-5 you had an unknown (say B) you would be able to solve for A and B you need a partial fraction expansion for this

OpenStudy (anonymous):

I've never seen something like this

OpenStudy (anonymous):

when do I will learn this?

OpenStudy (howard-wolowitz):

well If im correct this is part algebra and Trig

OpenStudy (solomonzelman):

yes, algebra

OpenStudy (anonymous):

so I think it will be next year

OpenStudy (triciaal):

4th degree polynomial 4 roots only 2 real showing

OpenStudy (solomonzelman):

\(\color{black}{ \displaystyle \frac{ A }{x-4} +\frac{ B }{x-5}=\frac{ -x+6 }{x^2-9x+20} }\) you see that (x-4)(x_5) is the same as \(x^2-9x+20\) but factored, (right?) multiply every term by the common denominator \(x^2-9x+20\), and you get: \(\color{black}{ \displaystyle \frac{ A }{x-4} +\frac{ B }{x-5}=\frac{ -x+6 }{x^2-9x+20} }\) \(\color{black}{ \displaystyle \frac{ A(x-5)(x-4) }{x-4} +\frac{ B(x-5)(x-4) }{x-5}=\frac{ (-x+6)(x-5)(x-4) }{x^2-9x+20} }\) \(\color{black}{ \displaystyle A(x-5) +B(x-4) =-x+6}\) plug in x=4 and find the A plug in x=5 and find the B

OpenStudy (solomonzelman):

in this case, you are already given the B, so you have \(\color{black}{ \displaystyle A(x-5) +1(x-4) =-x+6}\) \(\color{black}{ \displaystyle A(x-5) +x+4=-x+6}\) \(\color{black}{ \displaystyle A(x-5) +x=-x+2}\) \(\color{black}{ \displaystyle A(x-5) =-2x+2}\)

OpenStudy (solomonzelman):

it is still true for all x values, so you can plug any x value. try x=4

OpenStudy (triciaal):

I miss those days partial fractions are great

OpenStudy (howard-wolowitz):

that would give you 6

OpenStudy (solomonzelman):

I did it incorrectly

OpenStudy (solomonzelman):

A(x-5)+1(x-4)=-x+6 A(x-5)+x-4=-x+6 A(x-5)=-2x+10 now plug in x=6